# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

### v1

@@ Appendix C Detailed Results
Accuracy Similarity Validity Accuracy Similarity Validity Accuracy Similarity Validity
GPT-4o Achiam et al. (2023) 0.719 0.6586 0.8796 0.6864 0.642 0.8352 0.3952 0.618 0.857
GPT-4-turbo Achiam et al. (2023) 0.7662 0.6984 0.9048 0.7388 0.6821 0.8848 0.3946 0.6587 0.905
GPT-3.5-turbo Achiam et al. (2023) 0.4048 0.6327 0.854 0.412 0.6263 0.8486 0.3316 0.5635 0.8354
Claude-3.5 Anthropic (2024b) 0.797 0.7124 0.9422 0.6962 0.7112 0.911 0.5361 0.7042 0.8604
Claude-3 Anthropic (2024a) 0.7984 0.6067 0.9096 0.6094 0.6398 0.9062 0.4678 0.5855 0.9044
Gemini-1.5-pro Deepmind (2024) 0.7712 0.7022 0.9274 0.7876 0.6744 0.8926 0.4704 0.6077 0.9484
Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.5984 0.6028 0.6482 0.5684 0.6032 0.6272 0.2774 0.4828 0.634
Llama3-8B-Instruct Dubey et al. (2024) 0.4642 0.3658 0.6086 0.4332 0.4793 0.5704 0.2568 0.4547 0.6112
Llama3.1-8B-Instruct Dubey et al. (2024) 0.399 0.4235 0.5122 0.4164 0.483 0.5238 0.2655 0.4499 0.6158
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.222 0.4501 0.2802 0.1908 0.2578 0.3795 0.121 0.3244 0.2532
Qwen2-7B-Instruct Yang et al. (2024) 0 0.2923 0.0004 0.0002 0.4123 0.0004 0 0 0
Yi-1.5-9B 0.2884 0.5461 0.4927 0.205 0.3724 0.4126 0.1064 0.6596 0.4526
Chatglm-9B 0.3666 0.6902 0.4736 0.3514 0.682 0.5 0.1832 0.6506 0.4342
Llama-3.2-1B-Instruct Dubey et al. (2024) 0.0644 0.5055 0.1664 0.0822 0.441 0.1604 0.0714 0.4757 0.1796
MolT5-small 0.2158 0.1052 0.4302 0.2316 0.1011 0.442 0.2214 0.1031 0.4326
MolT5-base 0.2074 0.1051 0.4168 0.1856 0.1073 0.3796 0.2358 0.1054 0.4536
MolT5-large 0.4244 0.1015 0.8156 0.4496 0.1072 0.8678 0.4654 0.119 0.9214
BioT5-base 0.5158 0.1526 1 0.506 0.1597 1 0.5068 0.158 1
OpenMolIns-large (Llama-3.2-1B) 0.2898 0.5951 0.385 0.2644 0.5956 0.3678 0.1996 0.5849 0.349
OpenMolIns-large (Llama-3.1-8B) 0.8054 0.6678 0.872 0.7122 0.6548 0.8514 0.5224 0.6398 0.8802
OpenMolIns-light (Galactica-125M) 0.3202 0.6547 0.6416 0.3508 0.6435 0.6358 0.269 0.6521 0.638
OpenMolIns-small (Galactica-125M) 0.4172 0.642 0.5568 0.3958 0.6452 0.5338 0.2956 0.6385 0.5376
OpenMolIns-medium (Galactica-125M) 0.5904 0.5812 0.789 0.5874 0.5873 0.7384 0.4608 0.5859 0.7768
OpenMolIns-large (Galactica-125M) 0.6454 0.5927 0.8198 0.6388 0.5973 0.8028 0.495 0.5962 0.81
OpenMolIns-xlarge (Galactica-125M) 0.7362 0.5744 0.8902 0.7124 0.5697 0.8612 0.5786 0.5677 0.8626
Table 12: Results on MolOpt. For each task, we highlight the best accuracy and underline the second best accuracy.
@@
Accuracy Novelty Validity Accuracy Novelty Validity Accuracy Novelty Validity
GPT-4o Achiam et al. (2023) 0.1998 0.6703 0.5852 0.065 0.6336 0.8564 0.233 0.6513 0.859
GPT-4-turbo Achiam et al. (2023) 0.1702 0.6991 0.4904 0.0774 0.6301 0.9068 0.218 0.6605 0.8778
GPT-3.5-turbo Achiam et al. (2023) 0.107 0.5054 0.6947 0.0518 0.6871 0.5522 0.1136 0.6585 0.8686
Claude-3.5 Anthropic (2024b) 0.1928 0.6926 0.6548 0.1058 0.6584 0.886 0.2364 0.6582 0.8892
Claude-3 Anthropic (2024a) 0.1044 0.6833 0.591 0.1042 0.6598 0.8696 0.1816 0.9158 0.6644
Gemini-1.5-pro Deepmind (2024) 0.1742 0.6902 0.6774 0.0708 0.6522 0.8688 0.2486 0.6673 0.924
Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.1404 0.6675 0.5474 0.067 0.6478 0.7378 0.1752 0.6576 0.765
Llama3-8B-Instruct Dubey et al. (2024) 0.0242 0.6649 0.3812 0.026 0.6303 0.57 0.0848 0.6167 0.7216
Llama3.1-8B-Instruct Dubey et al. (2024) 0.0228 0.702 0.3862 0.0395 0.6541 0.6387 0.13 0.6274 0.6905
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.0078 0.6732 0.2986 0.0102 0.6309 0.4524 0.0048 0.6012 0.402
Qwen2-7B-Instruct Yang et al. (2024) 0.011 0.9061 0.2622 0.001 0.8645 0.0796 0.0022 0.8601 0.0622
Yi-1.5-9B Young et al. (2024) 0.0392 0.6848 0.617 0.0208 0.6407 0.7072 0.0126 0.6945 0.6521
Chatglm-9B GLM et al. (2024) 0.0002 0.7483 0.2131 0.0254 0.7189 0.4682 0 0.6908 0.5926
Llama-3.2-1B-Instruct Dubey et al. (2024) 0.004 0.6807 0.185 0.008 0.7465 0.2226 0.0008 0.7461 0.2818
MolT5-small Edwards et al. (2022) 0.0006 0.6586 0.661 0.0064 0.598 0.6202 0.0114 0.5287 0.8354
MolT5-base Edwards et al. (2022) 0.0008 0.6868 0.756 0.007 0.6509 0.8422 0.013 0.5464 0.8382
MolT5-large Edwards et al. (2022) 0.015 0.7103 0.8412 0.0118 0.5611 0.8916 0.0382 0.6088 0.9406
BioT5-base Pei et al. (2023) 0.0118 0.8353 0.995 0.0078 0.6667 0.9992 0.0476 0.6792 0.9998
OpenMolIns-large (LLama-3.2-1B) 0.0144 0.649 0.5616 0.035 0.615 0.6186 0.0252 0.6373 0.4412
OpenMolIns-large (LLama-3.1-8B) 0.0136 0.6634 0.7582 0.0544 0.6614 0.7456 0.1344 0.6396 0.6435
OpenMolIns-light (Galactica-125M) 0.0044 0.6054 0.793 0.0216 0.5724 0.7596 0.0244 0.5756 0.8442
OpenMolIns-small (Galactica-125M) 0.0146 0.6568 0.8424 0.053 0.6365 0.7926 0.057 0.5954 0.8874
OpenMolIns-medium (Galactica-125M) 0.0294 0.6553 0.8698 0.0622 0.6473 0.7474 0.0882 0.6091 0.8932
OpenMolIns-large (Galactica-125M) 0.0464 0.6729 0.9116 0.0716 0.6695 0.7374 0.0996 0.6276 0.8966
OpenMolIns-xlarge (Galactica-125M) 0.1862 0.6899 0.9308 0.1656 0.6887 0.7952 0.2006 0.6445 0.9162
Table 13: Results on MolCustom. For each task, we highlight the best accuracy and underline the second best accuracy.

### v2

@@ Appendix F Detailed Results
SR Similarity WSR Validity SR Similarity WSR Validity SR Similarity WSR Validity
GPT-4o (OpenAI, 2023) 0.7190 0.6586 0.4735 0.8796 0.6864 0.6420 0.4407 0.8352 0.3952 0.6180 0.2442 0.8570
GPT-4-turbo (OpenAI, 2023) 0.7662 0.6984 0.5351 0.9048 0.7388 0.6821 0.5039 0.8848 0.3946 0.6587 0.2599 0.9050
GPT-3.5-turbo (OpenAI, 2023) 0.4048 0.6327 0.2561 0.8540 0.4120 0.6263 0.2580 0.8486 0.3316 0.5635 0.1869 0.8354
Claude-3.5 (Anthropic, 2024b) 0.7970 0.7124 0.5678 0.9422 0.6962 0.7112 0.4951 0.9110 0.5361 0.7042 0.3775 0.8604
Claude-3 (Anthropic, 2024a) 0.7984 0.6067 0.4844 0.9096 0.6094 0.6398 0.3899 0.9062 0.4678 0.5855 0.2739 0.9044
Gemini-1.5-pro (Deepmind, 2024) 0.7712 0.7022 0.5415 0.9274 0.7876 0.6744 0.5312 0.8926 0.4704 0.6077 0.2859 0.9484
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.5984 0.6028 0.3607 0.6482 0.5684 0.6032 0.3429 0.6272 0.2774 0.4828 0.1339 0.6340
Llama3-8B-Instruct (Dubey et al., 2024) 0.4642 0.3658 0.1698 0.6086 0.4332 0.4793 0.2076 0.5704 0.2568 0.4547 0.1168 0.6112
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.3990 0.4235 0.1690 0.5122 0.4336 0.5257 0.2279 0.5910 0.2655 0.4499 0.1194 0.6158
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.2220 0.4501 0.0999 0.2802 0.1908 0.2578 0.0492 0.3795 0.1210 0.3244 0.0393 0.2532
Qwen2-7B-Instruct (Yang et al., 2024) 0.0000 0.2923 0.0000 0.0004 0.0002 0.4123 0.0001 0.0004 0.0000 0.0000 0.0000 0.0000
Yi-1.5-9B (Young et al., 2024) 0.2884 0.5461 0.1575 0.4927 0.2050 0.3724 0.0763 0.4126 0.1064 0.6596 0.0702 0.4526
Chatglm-9B (GLM et al., 2024) 0.3666 0.6902 0.2530 0.4736 0.3514 0.6820 0.2397 0.5000 0.1832 0.6506 0.1192 0.4342
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0644 0.5055 0.0326 0.1664 0.0822 0.4410 0.0363 0.1604 0.0714 0.4757 0.0340 0.1796
MolT5-small (Edwards et al., 2022) 0.2158 0.1052 0.0227 0.4302 0.2316 0.1011 0.0234 0.4420 0.2214 0.1031 0.0228 0.4326
MolT5-base (Edwards et al., 2022) 0.2074 0.1051 0.0218 0.4168 0.1856 0.1073 0.0199 0.3796 0.2358 0.1054 0.0249 0.4536
MolT5-large (Edwards et al., 2022) 0.4244 0.1015 0.0431 0.8156 0.4496 0.1072 0.0482 0.8678 0.4654 0.1190 0.0554 0.9214
BioT5-base (Pei et al., 2023) 0.5158 0.1526 0.0787 1.0000 0.5060 0.1597 0.0808 1.0000 0.5068 0.1580 0.0801 1.0000
Llama-3.2-1B (OpenMolIns-large) 0.2898 0.5951 0.1725 0.3850 0.2644 0.5956 0.1575 0.3678 0.1996 0.5849 0.1167 0.3490
Llama-3.1-8B (OpenMolIns-large) 0.8054 0.6678 0.5378 0.8720 0.7122 0.6548 0.4663 0.8514 0.5224 0.6398 0.3342 0.8802
Galactica-125M (OpenMolIns-light) 0.3202 0.6547 0.2096 0.6416 0.3508 0.6435 0.2257 0.6358 0.2690 0.6521 0.1754 0.6380
Galactica-125M (OpenMolIns-small) 0.4172 0.6420 0.2678 0.5568 0.3958 0.6452 0.2554 0.5338 0.2956 0.6385 0.1887 0.5376
Galactica-125M (OpenMolIns-medium) 0.5904 0.5812 0.3431 0.7890 0.5874 0.5873 0.3450 0.7384 0.4608 0.5859 0.2700 0.7768
Galactica-125M (OpenMolIns-large) 0.6454 0.5927 0.3825 0.8198 0.6388 0.5973 0.3816 0.8028 0.4950 0.5962 0.2951 0.8100
Galactica-125M (OpenMolIns-xlarge) 0.7362 0.5744 0.4229 0.8902 0.7124 0.5697 0.4059 0.8612 0.5786 0.5677 0.3285 0.8626
Table 14: Detailed results on MolOpt. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
@@
SR Novelty WSR Validity SR Novelty WSR Validity SR Novelty WSR Validity
GPT-4o (OpenAI, 2023) 0.1998 0.6703 0.1339 0.5852 0.0650 0.6336 0.0412 0.8564 0.2330 0.6513 0.1518 0.8590
GPT-4-turbo (OpenAI, 2023) 0.1702 0.6991 0.1190 0.4904 0.0774 0.6301 0.0488 0.9068 0.2180 0.6605 0.1440 0.8778
GPT-3.5-turbo (OpenAI, 2023) 0.1070 0.5054 0.0541 0.6947 0.0518 0.6871 0.0356 0.5522 0.1136 0.6585 0.0748 0.8686
Claude-3.5 (Anthropic, 2024b) 0.1928 0.6926 0.1335 0.6548 0.1058 0.6584 0.0697 0.8860 0.2364 0.6582 0.1556 0.8892
Claude-3 (Anthropic, 2024a) 0.1044 0.6833 0.0713 0.5910 0.1042 0.6598 0.0688 0.8696 0.1816 0.9158 0.1663 0.6644
Gemini-1.5-pro (Deepmind, 2024) 0.1742 0.6902 0.1202 0.6774 0.0708 0.6522 0.0462 0.8688 0.2486 0.6673 0.1659 0.9240
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.1404 0.6675 0.0937 0.5474 0.0670 0.6478 0.0434 0.7378 0.1752 0.6576 0.1152 0.7650
Llama3-8B-Instruct (Dubey et al., 2024) 0.0242 0.6649 0.0161 0.3812 0.0260 0.6303 0.0164 0.5700 0.0848 0.6167 0.0523 0.7216
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.0228 0.7020 0.0160 0.3862 0.0395 0.6541 0.0258 0.6387 0.1300 0.6274 0.0816 0.6905
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.0078 0.6732 0.0053 0.2986 0.0102 0.6309 0.0064 0.4524 0.0048 0.6012 0.0029 0.4020
Qwen2-7B-Instruct (Yang et al., 2024) 0.0110 0.9061 0.0100 0.2622 0.0010 0.8645 0.0090 0.0796 0.0022 0.8601 0.0019 0.0622
Yi-1.5-9B (Young et al., 2024) 0.0392 0.6848 0.0268 0.6170 0.0208 0.6407 0.0133 0.7072 0.0126 0.6945 0.0088 0.6521
Chatglm-9B (GLM et al., 2024) 0.0002 0.7483 0.0001 0.2131 0.0254 0.7189 0.0183 0.4682 0.0000 0.6908 0.0000 0.5926
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0040 0.6807 0.0027 0.1850 0.0080 0.7465 0.0060 0.2226 0.0008 0.7461 0.0006 0.2818
MolT5-small (Edwards et al., 2022) 0.0006 0.6586 0.0004 0.6610 0.0064 0.5980 0.0038 0.6202 0.0114 0.5287 0.0060 0.8354
MolT5-base (Edwards et al., 2022) 0.0008 0.6868 0.0005 0.7560 0.0070 0.6509 0.0046 0.8422 0.0130 0.5464 0.0071 0.8382
MolT5-large (Edwards et al., 2022) 0.0150 0.7103 0.0107 0.8412 0.0118 0.5611 0.0066 0.8916 0.0382 0.6088 0.0233 0.9406
BioT5-base (Pei et al., 2023) 0.0118 0.8353 0.0099 0.9950 0.0078 0.6667 0.0052 0.9992 0.0476 0.6792 0.0323 0.9998
Llama-3.2-1B (OpenMolIns-large) 0.0144 0.6490 0.0093 0.5616 0.0350 0.6150 0.0215 0.6186 0.0252 0.6373 0.0161 0.4412
Llama-3.1-8B (OpenMolIns-large) 0.0136 0.6634 0.0090 0.7582 0.0544 0.6614 0.0360 0.7456 0.1344 0.6396 0.0860 0.6435
Galactica-125M (OpenMolIns-light) 0.0044 0.6054 0.0026 0.7930 0.0216 0.5724 0.0124 0.7596 0.0244 0.5756 0.0140 0.8442
Galactica-125M (OpenMolIns-small) 0.0146 0.6568 0.0096 0.8424 0.0530 0.6365 0.0337 0.7926 0.0570 0.5954 0.0339 0.8874
Galactica-125M (OpenMolIns-medium) 0.0294 0.6553 0.0193 0.8698 0.0622 0.6473 0.0403 0.7474 0.0882 0.6091 0.0537 0.8932
Galactica-125M (OpenMolIns-large) 0.0464 0.6729 0.0312 0.9116 0.0716 0.6695 0.0479 0.7374 0.0996 0.6276 0.0625 0.8966
Galactica-125M (OpenMolIns-xlarge) 0.1862 0.6899 0.1285 0.9308 0.1656 0.6887 0.1140 0.7952 0.2006 0.6445 0.1293 0.9162
Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).